A straight Beam of ultrarelavistic electrons confined within a very long cylindrical region of radius less than 1 mm carries a current of 0.01 uA. The strength of the electric field at a point 2cm away from the beam axis will be nearest to
1
Expert's answer
2020-09-14T12:53:26-0400

As per the given question,

The radius of the cylindrical wire "(r)<1mm"

Current flowing in the wire "(I)=0.01 \mu A"

Electric field at a distance d=2cm from the wire =?


"B=\frac{\mu_o i}{2\pi r}"


But, "E=\frac{dB}{dt}"


So, "E=\frac{d}{dt}(\frac{\mu i}{2\pi r})"


"E=\frac{2\times 10^{-7}\times 0.01\times 10^{-6}}{2\times 10^{-2}}N/c"

"E=1\times 10^{-13} N/c"